HIGH-RESOLUTION ROTATIONAL SPECTROSCOPY OF THE CARBON CHAIN ANIONS C_{3}N^{−}, C_{4}H^{−}, AND C_{4}D^{−}

\underline{MICHAEL C. MCCARTHY} and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138.

The rotational spectra of C_{3}N^{−}, C_{4}H^{−}, and C_{4}D^{−} have been observed at high spectral resolution by Fourier transform microwave spectroscopy. For both C_{3}N^{−}and C_{4}D^{−}, frequencies of the hyperfine components in the lowest-J transitions have been determined to better than 0.1 ppm. The derived quadrapole coupling constants eQq for both anions are in good agreement with theoretical predictions. A number of other properties of these anions, including linewidths and rotational temperatures, have been systematically studied with respect to similar-sized neutral molecules. The production of C_{4}H^{−} using different hydrocarbon precursor and buffer gases has also been investigated.